OpenStudy (anonymous):

It takes a smaller hose twice as long to fill a small swimming pool as it does a larger hose. It takes both hoses working together 30 minutes to fill the swimming pool. How long will it take the smaller hose to fill the pool by itself?

OpenStudy (anonymous):

Let "s" be the time in minutes for the smaller hose to fill the swimming pool. Solve the following equation for s.\[\frac{1}{s}+\frac{1}{\frac{s}{2}}=\frac{1}{30} \]s = 90 minutes